callielovesyhuu:

Solve 4 over x minus 4 equals the quantity of x over x minus 4, minus four thirds for x and determine if the solution is extraneous or not.

Hero:

@Warriorz13 That's considered giving answers

Warriorz13:

._.

Warriorz13:

That's all ANYONE does, except maybe you.

Hero:

Do YOU, not what others do.

Warriorz13:

Yes I do.

Hero:

@callielovesyhuu, is this the correct setup? \[\dfrac{4}{x- 4} = \dfrac{x}{x - 4} - \dfrac{4}{3}\]

callielovesyhuu:

yes @Hero

Hero:

Okay, you can begin by subtracting \(\dfrac{x}{x - 4}\) from both sides. If you do so, you'll get: \[\dfrac{4}{x - 4} - \dfrac{x}{x - 4} = -\dfrac{4}{3}\]

Hero:

@callielovesyhuu what do you think is the next step here?

callielovesyhuu:

i got the answer now so thanks for your help

Hero:

Please show the work you did to arrive at the correct result.

Hero:

I'm curious to know how you arrived at the answer.

callielovesyhuu:

my teacher did it for me lol

Hero:

@callielovesyhuu show me the steps your teacher gave you.